Determine the LoD score for this RFLP marker to the gene associated with the trait assuming linkage of 10 map units (theta = 0.1) and then for 20 map units?   Does the value suggest linkage or non-linkage of the RFLP to the trait? Is the value significant?

DNA morph 2 morph 1 RE RE RE RE RE Probe P Probe P Southern Origin Morph 1 Morph 2 Heterozygote homozygote homozygote Linkage to D locus Dd dd 1 2 3 4 5 6 7 8 dd Dd Dd dd dd Dd Dd Dd Morph 1,2 2,2 2,2 1,2 1,2 2,2 2,2 1,2 1,2 2,2 Morph 1 Morph 2 RE Inference RE RERE RE Mother Crossover in child Father RE RE RE RE Morph 2 Morph 2
Answer #1

LOD = log10 likelihood if linked/ likelihood if the loci are unlinked

map unit = 10cm =10% recombinants

What is the probability of no recombinantion for each gamete?

1-r = prob(no recombination) = 0.90

there are two different parental genotypes

so, (1-r)/2 = 0.45

• Probability of 8 non-recombinants if genes are linked (at 10 cM)

We know , r=0.5 when loci are not linked

Probability if not linked is 0.25

Probability if not linked is 0.258

Lod score = log10 (0.458/0.258) = log10 (110.19) = 2.04              (The pattern of 8 non-recombinants is 110 times more likely if the genes are linked at 10 cM than if they were unlinked)

(LOD must be >3.0 to be “significant”)

So this is non significant.